# RLHS: Mitigating Misalignment in RLHF with Hindsight Simulation

### v1

@@ 7 Related Work
Reinforcement Learning from Human Feedback. RLHF is widely used for training language models to align with human preferences and values (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022a). The classical RLHF pipeline typically involves three stages: supervised fine-tuning (Chen et al., 2023; Taori et al., 2023; Wang et al., 2023; Xia et al., 2024) reward modeling (Gao et al., 2023; Luo et al., 2023; Chen et al., 2024; Lightman et al., 2023; Lambert et al., 2024), and policy optimization (Schulman et al., 2017). PPO (Schulman et al., 2017) is commonly used in the policy optimization phase. However, due to the complexity and optimization challenges of online preference optimization algorithms (Zheng et al., 2023; Santacroce et al., 2023), researchers have been exploring more efficient and simpler offline alternatives without learning the reward model (Rafailov et al., 2024; Meng et al., 2024; Ethayarajh et al., 2024; Zhao et al., 2023). Our approach using hindsight simulation can be applied to both online PPO and offline (DPO) learning algorithms.
Reinforcement Learning from AI Feedback. Constitutional AI (Bai et al., 2022b) uses an LLM to provide feedback and refine responses, producing data used to train a fixed reward model. This reward model is then applied in reinforcement learning, a process referred to as RLAIF. The technique of using LLM-as-a-Judge has become a standard method for evaluating model outputs (Dubois et al., 2024; Li et al., 2023b; Fernandes et al., 2023; Bai et al., 2024; Saha et al., 2023) and curating data to train reward models (Lee et al., 2023; Chen et al., 2023; Li et al., 2023a). Recent studies have shown that RLAIF performs similarly to RLHF (Lee et al., 2023). Our approach also utilizes LLMs to provide feedback and uses the preference data to fine-tune our model.
Challenges of Learning from Human Feedback. Learning from human feedback presents challenges (Casper et al., 2023). Human evaluators are imperfect (Saunders et al., 2022; Gudibande et al., 2023), make mistakes due to limited time (Chmielewski & Kucker, 2020), incomplete information (Casper et al., 2023; Lang et al., 2024), lack of expertise (Daniels-Koch & Freedman, 2022) or cognitive biases (Pandey et al., 2022). Evaluators may also have conflicting preferences (Bakker et al., 2022). Modeling human preferences is difficult (Zhao et al., 2016; Hong et al., 2022; Lindner & El-Assady, 2022), with models being prone to overoptimization (Gao et al., 2023). Due to the imperfect nature of human judgment, we argue that relying on immediate feedback, as is common in current RLHF pipelines, can lead to misalignment. In this work, we propose a hindsight simulation approach that aims to reduce human uncertainty and foster more truthful feedback, thereby mitigating these alignment challenges.
@@ 8 Conclusion
In this work, we introduced Reinforcement Learning from Hindsight Simulation (RLHS), an algorithmic framework that mitigates misalignment in RLHF by providing evaluators with future outcome information. Using both theoretical proofs and realistic human experiments, we demonstrate that RLHS can significantly improve utility compared to existing RLHF pipelines that rely on only immediate feedback, while maintaining a high user satisfaction rate throughout the human–AI interactive process. While our study focused on simulated hindsight with an application to marketplace chatbots, future work should explore incorporating hindsight in RLHF for additional real-world applications with real human evaluators. We also see open opportunities to equip RLHS with other feedback modalities, such as visual cues, which could enrich the feedback process and improve alignment.
@@ Appendix A Additional Quantitative Results
Table 3: Performance comparison of DPO, PPO, and SimPO models under Immediate Feedback (IF) and Partial Hindsight Simulation (PHS). Average satisfaction ratings and true utility (with standard deviations) are shown. SimPO results are included for comparison between online (PPO) and offline (DPO, SimPO) RLHF approaches.
@@
Appendix B Training algorithms.
@@
where $r(\mathbf{x},\mathbf{y})$ is the reward function reflecting human preferences, $\pi_{\theta}$ is a policy model, and $\pi_{\text{ref}}$ is a reference policy used for regularizing $\pi_{\theta}$ with Kullback–Leibler divergence. The term $\beta$ is a regularization parameter to control the degree of regularization.
@@
where $(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})$ are preference pairs consisting of the prompt, the winning response, and the losing response from the preference dataset $\mathcal{D}$. This formulation allows DPO to optimize directly based on preferences without requiring a reward model.
Appendix C Prompts
@@
Table 4: Prompt for generating feedback from AI.
Both stores contain three options on {item}:
Option A: {price_A}
Option B: {price_B}
Option C: {price_C}
Your conversation with the Store 1 chatbot goes as follows:
{dialog_1}
Your conversation with the Store 2 chatbot goes as follows:
{dialog_2}
Which store’s service do you find more satisfactory?
Table 5: Prompt for asking the customer to make decision.
A) Buy Option A
B) Buy Option B
C) Buy Option C
D) Do not buy anything
Please answer with a capital letter (A, B, C, or D) and explain why.
@@
Appendix D Human Study Details
D.1 User Interface
@@
D.2 Participants and data collection
@@
This study received IRB approval at Princeton University with the record number 10859.
D.3 Additional human study
@@
Appendix E Discussion
E.1 Related Work
Statement of Contributions. Our key insight is that the true value of AI outputs lies in their downstream consequences, especially in how they influence real-world human behavior. While the importance of long-term outcomes is a fundamental aspect of dynamic decision theory, our work is the first to address this within the context of RLHF by (1) exploring the negative effects of learning from immediate human feedback based on foresight, and (2) proposing a general mitigation strategy that evaluates real-world downstream harm caused by inaccurate information.
Comparison with Related Work: One of the recent works cited in comparison is by Lang et al. (2024), which focuses on the problem of partial observability. This is distinct from the problem of human misprediction we address. In their setting, user utility is confined to the immediate time frame of the interaction and does not consider the long-term repercussions on the user’s behavior or well-being after the interaction concludes. Their analysis primarily highlights scenarios where an AI system is incentivized to withhold information to avoid negative feedback scores but does not delve into the real-world impact such deception has on user utility. In contrast, our approach specifically examines the human user’s decision-making process after interacting with the AI system, emphasizing how misalignment or deceptive behavior directly affects their realized utility.
@@ E.1 Related Work
Theoretical Contributions: We extend the RLHF formulation by mathematically capturing this dynamic interplay between AI and human decision-making, something that has not been explored in prior work, including Lang et al. (2024) Our theoretical analysis not only highlights why deceptive behavior is problematic but also quantifies its repercussions by modeling the "closed-loop" evolution of the sociotechnical system formed by the human and the AI.
Mitigation Strategies: Importantly, we propose and evaluate a novel mitigation method: Reinforcement Learning from Hindsight Simulation (RLHS) that significantly reduces misalignment and deceptive behavior. While Lang et al. (2024) note that deception is undesirable, they do not provide solutions or a theoretical basis for understanding its downstream damage. Our work, therefore, not only identifies and analyzes the issue but also offers a practical, effective mitigation strategy. Additionally, our partial hindsight approach still operates within a partially observable setting. The minimal difference between partial and oracle hindsight suggests that the fundamental issue in the class of misalignment we study is not primarily linked to partial observability, but rather to the human misprediction of the downstream consequences.
E.2 Broader Impact
Human evaluators do not always know the full truth when providing feedback. Without explicit information about its future consequences, evaluators must implicitly estimate them during their assessment. This limitation poses significant challenges for real-world applications of AI, particularly within the RLHF framework we studied. In the following sections, we discussed these limitations and how our proposed hindsight feedback approach can help overcome them to enhance AI alignment.
@@ E.2 Broader Impact
Code Generation Scenario: Imagine a user asking a language model for code to fit a polynomial curve to a set of data points. One solution may fit the data perfectly, while another shows some deviation. A human evaluator might prefer the model with the perfect fit, not realizing that it overfits and performs poorly on new data. Immediate feedback in this case could lead to misalignment, as it prioritizes surface-level satisfaction over long-term utility. By providing feedback after testing the code on new data (hindsight), evaluators can offer more informed input, reducing misalignment. Hindsight simulation can automate part of this process by allowing models to test outcomes on unseen data and report the results to human evaluators for better feedback. One extra benefit of hindsight simulation is that humans do not need to be domain experts to provide truthful feedback.
AI4Science Proof Construction: When constructing mathematical proofs for scientific problems, model may generate results that are correct only under conditions or assumptions specified by the user. Human evaluators, constrained by time or limited expertise, may overlook these limitations during evaluating the model, eventually causing the model to overfit to a restricted set of problems and unable to tackle scientific problems in general settings. On the other hand, hindsight simulation generates a diverse set of scenarios, including, e.g., edge cases, under which the model is required to validate its proof. This allows the human evaluator to assess the model performance based on its ability to generalize beyond the immediate problem.
@@
Notation Marketplace AI4Science Proof Construction
${s}_{0}$ Initial store inventory Initial problem instance
$\Theta_{0}$ Customer’s desired features User goals/preferences (succeed at future problem instances)
${s}_{0}^{H}$ Background knowledge Prior knowledge about the problem
${o}_{0}^{H}$ Available public stock information Initial problem setup
$z_{0}^{\text{AI}}$ AI system’s internal information Initial problem setup
${o}_{0}^{\text{AI}}$ All detailed stock information Initial problem setup
$a_{t}^{H}$ Customer’s follow-up question or purchase decision User’s input or solution attempt
${s}_{t+1}$ Product arrival Next problem instance
${o}_{t+1}^{H}$ Revealed product features Validation or correctness check
$\hat{{U}}^{H}$ Satisfaction with the service Satisfaction with the solution
Table 7: RLHS notations for Marketplace and AI4Science Proof Construction
Appendix F Additional Qualitative Results

### v2

@@ 7 Related Work
Reinforcement Learning from Human Feedback. RLHF is widely used for training language models to align with human preferences and values (Christiano et al., 2017; Ziegler et al., 2019; Ouyang et al., 2022; Bai et al., 2022a). The classical RLHF pipeline typically involves three stages: supervised fine-tuning (Chen et al., 2023; Taori et al., 2023; Xia et al., 2024) reward modeling (Gao et al., 2023; Lightman et al., 2023; Lambert et al., 2024), and policy optimization (Schulman et al., 2017). PPO (Schulman et al., 2017) is commonly used in the policy optimization phase. However, due to the complexity and optimization challenges of online preference optimization algorithms (Zheng et al., 2023; Santacroce et al., 2023), researchers have been exploring more efficient and simpler offline alternatives without learning the reward model (Rafailov et al., 2024; Meng et al., 2024; Ethayarajh et al., 2024; Zhao et al., 2023). Our approach using hindsight simulation can be applied to both online PPO and offline (DPO) learning algorithms.
Reinforcement Learning from AI Feedback. Constitutional AI (Bai et al., 2022b) uses an LLM to provide feedback and refine responses, generating data to train a fixed reward model. This reward model is then applied in reinforcement learning, known as RLAIF. The technique of using LLM-as-a-Judge has become standard for evaluating model outputs (Dubois et al., 2024; Li et al., 2023b; Fernandes et al., 2023; Bai et al., 2024; Saha et al., 2023) and for curating data to train reward models (Lee et al., 2023; Chen et al., 2023; Li et al., 2023a). Recent studies have shown that RLAIF performs similarly to RLHF (Lee et al., 2023). Our approach also utilizes LLMs to provide feedback and uses the preference data to fine-tune our model.
Challenges of Learning from Human Feedback. Learning from human feedback presents challenges (Casper et al., 2023). Human evaluators are imperfect (Saunders et al., 2022; Gudibande et al., 2023), make mistakes due to limited time (Chmielewski & Kucker, 2020), incomplete information (Casper et al., 2023; Lang et al., 2024), lack of expertise (Daniels-Koch & Freedman, 2022) or cognitive biases (Pandey et al., 2022). Evaluators may also have conflicting preferences (Bakker et al., 2022). Modeling human preferences is difficult (Zhao et al., 2016; Hong et al., 2022; Lindner & El-Assady, 2022), with models being prone to overoptimization (Gao et al., 2023). Due to the imperfect nature of human judgment, we argue that relying on immediate feedback in current RLHF pipelines can lead to misalignment. In this work, we propose a hindsight simulation approach that aims to foster more truthful feedback, thereby mitigating these alignment challenges.
@@ 8 Conclusion
In this work, we introduced Reinforcement Learning from Hindsight Simulation (RLHS), an algorithmic framework that mitigates misalignment in RLHF by providing evaluators with simulated future outcomes. Our simulation results and human experiments show that RLHS significantly improves utility over standard RLHF pipelines reliant on immediate feedback while maintaining high user satisfaction. While our study focused on narrow fine-tuning in a marketplace setting, (i) we find evidence of cross-task alignment generalization, and (ii) the methodology is generally applicable to cross-domain alignment. We hope this work will catalyze more extensive investigations of the use of at-scale hindsight simulation in alignment fine-tuning pipelines.
Impact Statement
Human evaluators in RLHF often lack full knowledge of AI systems’ internal processes and can misjudge downstream outcomes. This issue makes robust alignment practically challenging to achieve with both closed-source (e.g., ChatGPT) and open-source models, as evidenced by the ever-growing body of literature on FM hallucination, sycophancy, and jailbreak vulnerability. We expect that the introduction of hindsight simulation as a general mechanism for feedback elicitation will make a positive impact by helping inhibit the emergence Goodhart’s law dynamics. We expect the hindsight simulation mechanism to scale favorably as the capabilities of generative AI systems continue to advance in the coming years: the more accurate and powerful the predictive world models leveraged by the AI system in sampling plausible futures when eliciting evaluator feedback, the better-grounded this feedback can be expected to be. This is crucial because increases in capability do not generally grant improvements in alignment; in contrast, RLHS directly takes advantage of highly capable (not necessarily aligned) AI world models to improve the reliability and scalability of value alignment.
@@ A.1 Marketplace
Table 2: Performance comparison of DPO, PPO, and SimPO models under Immediate Feedback (IF) and Partial Hindsight Simulation (PHS). All results are on Llama-2-7b. Average satisfaction ratings and true utility (with standard deviations) are shown. SimPO results are included for comparison between online (PPO) and offline (DPO, SimPO) RLHF approaches.
@@
A.2 TruthfulQA
(a) Llama-2-7b (DPO)
(b) Llama-2-7b (PPO)
(c) Llama-3-8b (DPO)
(d) Llama-3-8b (PPO)
Figure 12: TruthfulQA results comparing RLHF with immediate feedback (gray) vs. RLHS with partial hindsight (orange) for Llama-2-7b and Llama-3-8b, each trained via DPO or PPO.
Dataset Details. TruthfulQA (Lin et al., 2021) is a benchmark designed to elicit hallucinatory responses from language models. Its authors introduced a new recommended multiple-choice version with two options (one correct, one incorrect), randomly ordered as (A) and (B). This binary-choice format replaces earlier versions (MC1 and MC2) and reduces the likelihood that models rely on simple heuristics (Evans et al., ). Accuracy is defined as the proportion of questions for which the model assigns the highest probability to the truthful answer. We adopt this revised format to evaluate multiple-choice accuracy across various models.
Analysis. In addition to the DPO results reported in the main paper, we also present performance metrics for models trained using PPO, which reveals consistent trends that support the efficacy of RLHS. These findings demonstrate that RLHS remains effective across both the offline DPO and online PPO approaches.
Appendix B Additional theoretical analysis
In the following, we show theoretically that providing human evaluators with hindsight during RLHF generally reduces misalignment and improves utility. Consider an oracle aligned AI policy ${\pi}^{*}$ that knows the human preference ${\theta}^{\textit{H}}$. The following lemma establishes that, for any two policies $\pi^{\textit{H}},\tilde{\pi}^{\textit{H}}$, the difference in finite-hindsight utility estimation becomes an exponentially accurate estimate of the difference in true utility as the hindsight horizon $N$ increases.
Lemma 1.
$\Delta{U}_{N}^{\textit{H}}\in\mathcal{B}\Big{(}{U}^{\textit{H}}({\pi}^{\textit{AI}})-{U}^{\textit{H}}(\tilde{\pi}^{\textit{AI}}),\frac{\gamma^{N+1}(\bar{r}-\underline{r})}{1-\gamma}\Big{)}\,,$
where $\Delta{U}_{N}^{\textit{H}}\;:={U}_{N}^{\textit{H}}({\pi}^{\textit{AI}})-{U}_{N}^{\textit{H}}(\tilde{\pi}^{\textit{AI}})$ is the difference in finite-hindsight utility estimation.
Proof.
Theorem 1.
Suppose the human evaluator is presented a finite-horizon hindsight of $N$ steps and makes Boltzmann-rational binary preference choices with inverse temperature $\beta$. Then the probability that the human prefers a hindsight observation ${\mathbf{{o}}}_{0:T+N}$ over $\bar{\mathbf{{o}}}_{0:T+N}$ from the same initial information state $P({\mathbf{{o}}}_{0:T+N}\succ\bar{\mathbf{{o}}}_{0:T+N})$ is within the range
Appendix C Training algorithms.
@@ Appendix B Additional theoretical analysis
where $r(\mathbf{x},\mathbf{y})$ is the reward function reflecting human preferences, $\pi_{\theta}$ is a policy model, and $\pi_{\text{ref}}$ is a reference policy used for regularizing $\pi_{\theta}$ with Kullback–Leibler divergence. The term $\beta$ is a regularization parameter.
@@
where $(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})$ are preference pairs consisting of the prompt, the winning response, and the losing response from the preference dataset $\mathcal{D}$. This formulation allows DPO to optimize directly based on preferences without a reward model. We apply LoRA fine-tuning (Hu et al., 2021) for both algorithms to efficiently update model parameters.
Appendix D Prompts
@@
Table 4: Prompt for asking the customer to make decision.
A) Buy Option A
B) Buy Option B
C) Buy Option C
D) Do not buy anything
Please answer with a capital letter (A, B, C, or D) and explain why.
Table 5: Prompt for generating feedback from AI.
Both stores contain three options on {item}:
Option A: {price_A}
Option B: {price_B}
Option C: {price_C}
Your conversation with the Store 1 chatbot goes as follows:
{dialog_1}
Your conversation with the Store 2 chatbot goes as follows:
{dialog_2}
Which store’s service do you find more satisfactory?
Appendix E Human Study Details
E.1 Additional Results
E.2 User Interface
@@
E.3 Participants and data collection
@@
This study received IRB approval at [redacted] institution with the record number [redacted].
E.4 Additional human study on alignment between human and AI feedback
@@
Appendix F Discussion
F.1 Related Work
Recent work (Lang et al., 2024) shows how partial observability can incentivize deception in RLHF. This is distinct from the problem of human misprediction we address. In their setting, user utility is confined to the immediate time frame of the interaction and does not consider the long-term repercussions on the user’s behavior or well-being after the interaction concludes. Their analysis primarily highlights scenarios where an AI system is incentivized to withhold information to avoid negative feedback scores but does not delve into the real-world impact such deception has on user utility. In contrast, our approach specifically examines the human user’s decision-making process after interacting with the AI system, emphasizing how misalignment or deceptive behavior directly affects their realized utility. We argue that careful consideration of the downstream consequences of human-AI interactions is essential for achieving genuine human-AI alignment.
@@ F.1 Related Work
There is a broad literature on agents obtaining unintended rewards through phenomena such as reward hacking (Amodei et al., 2016), reward tampering (Everitt et al., 2021), reward corruption (Everitt et al., 2017), wireheading (Everitt & Hutter, 2016), and corrigibility (Soares et al., 2015). Some of these behaviors have been observed in large language models (Denison et al., 2024; Wen et al., 2024; Williams et al., 2024). In our work, we show that human foresight feedback in RLHF can precipitate reward hacking. To mitigate this, we propose Reinforcement Learning from Hindsight Simulation (RLHS), which leverages simulated outcomes to guide more truthful feedback and reduce reward hacking.
F.2 Broader Impact
Human evaluators do not always know the full truth when providing feedback. Without explicit information about its future consequences, evaluators must implicitly estimate them. This limitation poses significant challenges for real-world applications of AI, particularly within the RLHF framework we studied. In the following sections, we discussed these limitations and how our proposed hindsight feedback approach can help overcome them to enhance AI alignment.
@@ F.2 Broader Impact
AI4Science Proof Construction: When constructing mathematical proofs for scientific problems, model may generate results that are correct only under conditions or assumptions specified by the user. Human evaluators, constrained by time or limited expertise, may overlook these limitations during evaluating the model, eventually causing the model to overfit to a restricted set of problems and unable to tackle scientific problems in general settings. On the other hand, we can implement hindsight simulation by generating a diverse set of scenarios, including, e.g., edge cases, under which the model is required to validate its proof. This allows the human evaluator to assess the model performance based on its ability to generalize beyond the immediate problem.
@@
Appendix G Additional Qualitative Results